How to solve this question please given the solution fast it`s urgentSolve sin2x = cos3x

Grade:11

1 Answers

Laltesh Singh
6 years ago
Sin2x = cos3xCos (π/2 - 2x) = cos3x2nπ + π/2 -2x = 3x2nπ + π/2 = 5xπ (2n + 1/2) = 5xx = π/5 (2n + 1/2)